OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the given point. (4x^4)+(4x^2)-8x with point of (1,0)

OpenStudy (anonymous):

\[4x ^{4}+4x ^{2}-8x\]

myininaya (myininaya):

y'=16x^3+8x-8=8(2x^3+x-1) y' at x=1 is 8(2+1-1)=16 y' represents the slope so the slope is 16 y=mx+b y=16x+b we know a point on the tangent line (1,0) 0=16(1)+b so we can solve for b so b=-16 so the tangent line is y=16x-16

OpenStudy (anonymous):

Why you use x =1? when you solved to get 16

myininaya (myininaya):

we want to find the tangent line at x=1 so we need to find the slope at x=1

OpenStudy (anonymous):

so it it possinle yo use any other x value?

OpenStudy (anonymous):

Or becasue the opint given is (1,0)

myininaya (myininaya):

it depends at what point they want the tangent line they asked for the tangent line at x=1 so we found the tangent line at x=1

OpenStudy (anonymous):

ok well now i get thanks again

myininaya (myininaya):

(x=1,y=0)